Let

T:

R3→R2

be a linear transformation such that


(1,1,1) = (1,0),


(1,1,0) = (2, −1) and


(1,0,0) =


(4,3). What is


(2, −3,5)?


1
Expert's answer
2021-11-15T15:19:36-0500
"T\bigg(\begin{bmatrix}
 1 \\
 1 \\
1
\end{bmatrix}\bigg)=\begin{bmatrix}
 1 \\
 0
\end{bmatrix},"

"T\bigg(\begin{bmatrix}
 1 \\
 1 \\
0
\end{bmatrix}\bigg)=\begin{bmatrix}
 2 \\
 -1
\end{bmatrix},"

"T\bigg(\begin{bmatrix}
 1 \\
 0 \\
0
\end{bmatrix}\bigg)=\begin{bmatrix}
 6 \\
 3
\end{bmatrix}"

"A=\begin{bmatrix}
 a & b & c \\
 d & e & f
\end{bmatrix}"

"\begin{matrix}
 a+b+c=1 \\
 d+ e+f=0
\end{matrix}"

"\begin{matrix}
 a+b=2 \\
 d+ e=-1
\end{matrix}"

"\begin{matrix}
 a=4 \\
 d=3
\end{matrix}"

"A=\begin{bmatrix}
 4 & -2 & -1 \\
 3 & -4 & 1
\end{bmatrix}"

"\begin{bmatrix}
 4 & -2 & -1 \\
 3 & -4 & 1
\end{bmatrix}\begin{bmatrix}
 2 \\
 -3 \\
5
\end{bmatrix}=\begin{bmatrix}
 4(2)-2(-3)-1(5) \\
 3(2)-4(-3)+1(5) 
\end{bmatrix}"

"=\begin{bmatrix}
 9 \\
 23
\end{bmatrix}"